# Spinning masters: on the impact of tidal forces and protohalo size on early spin evolution

### v1

@@ Abstract
In this work, we explore how the size and surrounding tidal fields of dark matter protohalos at high redshift influence their angular momentum (AM) evolution. While tidal torque theory (TTT) states that AM arises from the misalignment between protohalo shape and tidal fields, it remains unclear what is the characteristic scale of the perturbations that couple with each protohalo, and its correlation with protohalo properties such as size. Moreover, although the assumptions of the TTT are assumed to hold during the linear and quasi-linear regime, cosmological simulations reveal that discrepancies between its predictions and the true AM of halos emerge earlier than expected. To address this, we analyze cosmological simulations to study tidal fields at $z=80$ using different smoothing lengths, and determine which best predicts AM under TTT. We then investigate discrepancies between predicted and actual AM across redshifts, considering the effect of evolving tidal and inertia tensors. Our results show that the early tidal field couples with the inertia tensor of protohalos on scales about half of their characteristic size and confirm that disagreements between theory and simulation arise before the non-linear regime, suggesting a systematic effect from protohalo shape interacting with the forming cosmic web.
@@
methods: numerical – methods: statistical – large-scale structure of Universe – galaxies: haloes
@@ 1 Introduction
The acquisition of angular momentum (AM) by galaxies and dark matter (DM) halos, first studied by Hoyle et al. (1949), is crucial for galaxy formation and its link to large-scale structure (LSS). AM growth and orientation impact halo-galaxy models (Zentner et al., 2014; Rodriguez et al., 2015; McEwen & Weinberg, 2018) and introduce weak lensing systematics via intrinsic alignments (Troxel & Ishak, 2015; Hikage et al., 2019; Fabbian et al., 2019; Copeland et al., 2020). The spin parameter affects galactic disk properties (Peebles, 1969) and drives halo assembly bias (Sheth & Tormen, 2004; Gao et al., 2005; Wechsler et al., 2006; Bett et al., 2007; Gao & White, 2007), though its origins remain unclear (Montero-Dorta et al., 2024), including potential cosmological dependencies (Lazeyras et al., 2021; Contreras et al., 2021). AM-LSS alignments may also constrain neutrino mass (Lee et al., 2020).
In hierarchical structure formation, DM halos arise from collapsing primordial fluctuations, creating the gravitational framework for galaxy evolution. These structures assemble into the cosmic web, a network of nodes, filaments, walls, and voids shaped significantly by tidal fields (Bond et al., 1996; van de Weygaert & Bond, 2008; Aragón-Calvo et al., 2010; Cautun et al., 2014; Feldbrugge & van de Weygaert, 2024; Kugel & van de Weygaert, 2024; Libeskind et al., 2018). The same tidal forces also transfer AM to halos and galaxies during collapse.
@@
While TTT has been successful in explaining the initial acquisition of AM and its early alignment with the LSS in a statistical sense, several critical aspects remain poorly understood. In this work, we focus on unresolved questions, such as the role of different tidal field scales in AM acquisition, the connection between these scales and protohalo properties, and the timing and nature of deviations from TTT predictions during protohalo evolution. By addressing these gaps, we aim to advance our understanding of AM evolution, its connection to present-day halo properties and, if possible, with observable galaxy properties.
@@
Numerical simulations have played a crucial role in validating and refining TTT. While the theory accurately describes AM growth in the linear regime (White, 1984), simulations reveal significant deviations during non-linear evolution (Sugerman et al., 2000; Lee & Pen, 2000; Porciani et al., 2002a, b). These discrepancies emerge as environmental effects, mergers, and vortical flows become important (Vitvitska et al., 2002; Libeskind et al., 2013; Laigle et al., 2015; Bett & Frenk, 2016; López et al., 2019; Lee & Moon, 2022). Such results highlight the need to extend TTT with additional physics that accounts for non-linear processes and environmental factors (e.g. Codis et al., 2015; Wang & Kang, 2017; Neyrinck et al., 2020; Ebrahimian & Abolhasani, 2021; Moon & Lee, 2023; López & Merchán, 2023).
@@
First, while TTT provides an elegant framework, simulations reveal significant deviations in DM halo AM evolution from its predictions. This suggests either theoretical incompleteness or gaps in our understanding of certain aspects – particularly the coupling scale between tidal fields and protohalo inertia tensors in Lagrangian space. While studies show that varying density field smoothing scales affect collapse times and inertia-tidal alignments (Lee et al., 2009; Hahn et al., 2009; Moon & Lee, 2024), the conventional assumption that protohalo size determines the relevant TTT scale lacks rigorous validation. Critical unanswered questions remain: How do tidal field properties systematically vary with filtering scale? Which scales dominate tidal-inertia tensor correlations? How are these scales linked to protohalo characteristics? A thorough analysis of these scale-dependent effects is notably absent in current literature.
@@ 2 Brief outline of the TTT
Figure 1: Five examples of the evolution of protohalos, with increasing mass from top to bottom. Each column, from left to right, shows the particle distribution at $z=80,4.5,1.5,0.5$ and 00, respectively. The black dots represent the full FoF group, while the red dots show the particles that reside within $0.1R_{\mathrm{vir}}$ at present time (see Section 6). Notice that the axes that correspond to each protohalo are scaled differently.
This equation indicates that, during the early stages of structure formation, the AM arises from the misalignment between the shape of the protohalo and the principal directions of expansion or collapse associated to its surrounding tidal field appropriately smoothed. The ambiguity of this expression is deliberate: in general, it is not prescribed what is the scale of the tidal field that couples with the inertia tensor to produce AM. It is usually assumed that corresponds roughly to the size of each protohalo. However, (1) the boundaries of a protohalo, and hence its size and shape, are poorly defined in Lagrangian space, and (2) the amplitude and orientation of the tidal field can change with the smoothing scale both below and above the characteristic size of each protohalo. Hence, even if we could define the size of a protohalo, $R_{\mathrm{h}}$, unequivocally, the tidal field and hence the TTT predictions would be different when considering a density field smoothed over e.g. $R_{\mathrm{h}}/2$, $R_{\mathrm{h}}$ or $2R_{\mathrm{h}}$.
@@ 3.1 The simulation, halos and protohalos
In this work we use the code gadget 2 (Springel, 2005) to simulate the cosmological evolution of $1600^{3}$ dark matter particles in a periodic cube of side 400 $400{\textrm{ $h^{-1}$Mpc}}$, with a mass per particle $m_{\mathrm{p}}=1.18219\times 10^{9}{\textrm{ $h^{-1}$M${}_{\sun}$}}$. This simulation consists of 205 outputs between $z=80$, when the initial conditions were generated, and $z=0$, which corresponds to the present time. Cosmological parameters were taken from the Planck Collaboration results (Planck Collaboration et al., 2020): matter density $\Omega_{\mathrm{m}}=1-\Omega_{\Lambda}=0.315$, Hubble constant $H_{0}=67.4{\textrm{ km\,s${}^{-1}$}}{\textrm{ Mpc}}^{-1}$ and normalization parameter $\sigma_{8}=0.811$. Dark matter halo identification was performed at $z=0$ using a standard Friends of Friends (FoF) algorithm with percolation length $l=0.17\bar{\nu}^{-1/3}$, where $\bar{\nu}$ is the mean particle number density. In order to avoid biases in the determination of dynamical properties due to the low number of particles (see e.g. Paz et al., 2006; Bett et al., 2007), our analysis was carried out discarding FoF groups with less than $250$ particles. The result is a sample of about $6.5\times 10^{5}$ halos with masses $_{\mathrm{FoF}}\geqslant 3\times 10^{11}{\textrm{ $h^{-1}$M${}_{\sun}$}}$ M.
@@
In addition to analyzing the full FoF groups, we track the particles that, by $z=0$, reside within fixed fractions of the virial radius ($0.1R_{\mathrm{vir}}$ and $0.5R_{\mathrm{vir}}$) of each halo (see Section 6). By comparing these inner particles to their outer counterparts, we aim to understand whether the AM assembly is driven differentially (e.g., if inner particles experience distinct tidal torques) or non-linear effects) or if the entire Lagrangian region evolves coherently. This distinction is important for understanding how the AM of the final halo relates to its initial conditions.
Figure 1 presents the Lagrangian evolution of five representative FoF groups (black dots) and their inner $0.1R_{\mathrm{vir}}$ particles (red dots) at redshifts $z=80,4.5,1.5,0.5$, and 00. The full FoF groups exhibit diverse initial configurations, ranging from contiguous to fragmented spatial distributions, particularly for higher-mass halos. This morphological variety demonstrates the complexity of characterizing Lagrangian regions. The inner particles show similarly irregular distributions, either scattered throughout the entire volume or partially clustered in subregions.
The evolutionary sequence reveals distinct phases of collapse. By $z=4.5$, particles remain broadly distributed, with only faint filamentary structures emerging (a pattern reminiscent of large-scale structure formation). The distribution becomes increasingly clumpy by $z=1.5$, with most inner particles forming an offset core that becomes prominent by $z=0.5$.
Figure 2: Protohalo size as a function of their mass. The dot, solid and dashed lines show, respectively, the median of the equivalent radius ($R_{\mathrm{eq}}$), the Lagrangian radius ($R_{\mathrm{lag}}$), and ten times the virial radius ($10R_{\mathrm{vir}}$) for protohalos of different mass. Although the value of $R_{\mathrm{lag}}$ is computed exactly from the halo mass, $R_{\mathrm{eq}}$ and $10R_{\mathrm{vir}}$ present some scatter which we show as the dispersion of the gray circles and crosses. The colored regions indicate the three mass bins we use in our analysis.
@@
Finally, we need to characterize the large-scale structure at present time. To this end, we implement the code NEXUS+ (Cautun et al., 2013) at $z=0$, which establishes the type of region in which each halo is embedded through an analysis of the Hessian of the smoothed local density field at different scales. In this way, it identifies not only the cosmological environment (void, wall, filament or node), but also the preferred directions of collapse and expansion of matter, $\hat{\mathbf{e}}_{k}$, with $k=1,2,3$. In the case of filamentary regions, for example, $\hat{\mathbf{e}}_{3}$ can be associated with the direction of the main axis or spine. A more detailed description of NEXUS+ can be found on Appendix A.
@@ 3.1.1 The size of protohalos
Our second measure of protohalo size is the equivalent radius, which we define as:
$R_{\mathrm{eq}}=(abc)^{1/3},$
@@
where $a,b,c=\sqrt{\mathrm{i}_{k}}$ for $k=1,2,3$. This is simply the radius of a sphere equivalent in volume to the ellipsoid associated with the eigenvalues of the inertia tensor. This quantity is computed using the particle distribution of each protohalo at $z=80$, i.e., at the same time when we analyze its surrounding tidal field.
Finally, we have $10R_{\mathrm{vir}}$, i.e. ten times the virial radius of each halo at $z=0$. The factor multiplying $R_{\mathrm{vir}}$ aims to take into account that, in comoving coordinates, the particle distribution of each protohalo at $z=80$ is significantly extended with respect to the virialized object at $z=0$. The value of this factor has been chosen such that $10R_{\mathrm{vir}}$ represents a scale typically larger than the protohalo itself. Moon & Lee (2024) found that calculating the tidal field using a similar filter length, $8R_{\mathrm{vir}}$, yields a universal correlation with the inertia tensor, independent of the mass of each protohalo. Moreover, they found that $8R_{\mathrm{vir}}$ establishes an upper limit beyond which the so-called “type II spin transition” is not observed, i.e., the transition of $\hat{\mathbf{J}}$ from $\perp\hat{\mathbf{t}}_{1}$ to $\parallel\hat{\mathbf{t}}_{1}$, where $\hat{\mathbf{t}}_{1}$ represents the first axis of the tidal field. Thus, the value $10R_{\mathrm{vir}}$ allows us to test these results and compare our conclusions.
@@
Figure 2 shows the relation between halo mass and the three different measures of the protohalo size at $z=80$: the equivalent radius ($R_{\mathrm{eq}}$), the Lagrangian radius ($R_{\mathrm{lag}}$) and ten times the virial radius ($10R_{\mathrm{vir}}$). Notice that, whereas the value of $R_{\mathrm{lag}}$ is computed exactly from the halo mass, $R_{\mathrm{eq}}$ and $10R_{\mathrm{vir}}$ present some scatter which can be seen as the dispersion of the gray circles and crosses, respectively. Naturally, the protohalo size tends to increase with halo mass. The equivalent radius spans a range of values such that its median goes from $\sim 0.5$ to 2.6 $2.6{\textrm{ $h^{-1}$Mpc}}$ in the mass range $[3\times 10^{11}-5\times 10^{13}]{\textrm{ $h^{-1}$M${}_{\sun}$}}$. Typically, for each protohalo, $R_{\mathrm{eq}}$ is about half of the value of $R_{\mathrm{lag}}$ and $\sim 1/3$ of $10R_{\mathrm{vir}}$.
@@
In order to consider the possible dependence of our results with halo mass, we split our halo population in three subsamples, shown in Figure 2 as coloured regions. With $m=\log\left[M_{\mathrm{FoF}}/({\textrm{ $h^{-1}$M${}_{\sun}$}})\right]$, we define a low mass subsample within the range $m=[11.5,12.0]$, a subsample of intermediate mass $m=[12.0,12.7]$ and, finally, a high mass subsample with masses in the range $m=[12.7,13.7]$. These ranges have been chosen such that a well known trend in spin alignment at $z=0$ can be easily identified: we expect the spin of the low (high) mass subsample to be preferentially aligned (perpendicular) to the present day filaments.

### v2

@@ Abstract
In this work, we explore how the size and surrounding tidal fields of dark matter protohalos at high redshift influence their angular momentum (AM) evolution. While tidal torque theory (TTT) states that AM arises from the misalignment between protohalo shape and tidal fields, it remains unclear what is the characteristic scale of the perturbations that couple with each protohalo, and its correlation with protohalo properties such as size. Moreover, although the assumptions of the TTT are assumed to hold during the linear and quasi-linear regime, cosmological simulations reveal that discrepancies between its predictions and the true AM of halos emerge earlier than expected. To address this, we analyze cosmological simulations to study tidal fields at $z=80$ using different smoothing lengths, and determine which best predicts AM under TTT. We then investigate discrepancies between predicted and actual AM across redshifts, considering the effect of evolving tidal and inertia tensors. Our results show that the early tidal field couples with the inertia tensor of protohalos on scales about half of their characteristic size and confirm that disagreements between theory and simulation emerge already at relatively early cosmic times ($z\sim 10-5$), suggesting a systematic effect from protohalo shape interacting with the forming cosmic web.
@@
methods: numerical – methods: statistical – large-scale structure of Universe – galaxies: halos
@@ 1 Introduction
The acquisition of angular momentum (AM) by galaxies and dark matter (DM) halos, first studied by Hoyle et al. (1949), is crucial for galaxy formation and its link to large-scale structure (LSS). AM growth and orientation impact halo-galaxy models (Zentner et al., 2014; Rodriguez et al., 2015; McEwen & Weinberg, 2018) and introduce weak lensing systematics via intrinsic alignments (Troxel & Ishak, 2015; Hikage et al., 2019; Fabbian et al., 2019; Copeland et al., 2020). Galaxy formation models often assume a close coupling between dark matter and baryons and conservation of angular momentum, leading to a natural link between disk size and halo size through the halo spin parameter (Peebles, 1971; Fall, 1979; Dalcanton et al., 1997; Mo et al., 1998; Somerville et al., 2008; Firmani & Avila-Reese, 2009; Dutton et al., 2011). The dependence of clustering on halo spin can be regarded as one manifestation of assembly bias (Sheth & Tormen, 2004; Gao et al., 2005; Wechsler et al., 2006; Bett et al., 2007; Gao & White, 2007), though the origin of this correlation remains unclear (Montero-Dorta et al., 2024), including possible cosmological dependencies (Lazeyras et al., 2021; Contreras et al., 2021). AM alignments with respect to the large-scale structure may also place constraints on neutrino mass (Lee et al., 2020).
In hierarchical structure formation, DM halos arise from collapsing primordial fluctuations, creating the gravitational framework for galaxy evolution. These structures assemble into the cosmic web, a network of nodes, filaments, walls, and voids shaped significantly by tidal fields (e.g. Zel’dovich, 1970; Einasto et al., 1977; Shandarin & Zeldovich, 1989; Bond et al., 1996; Hahn et al., 2007; van de Weygaert & Bond, 2008; Lee et al., 2009; Aragón-Calvo et al., 2010; Hahn et al., 2010; Cautun et al., 2014; Feldbrugge & van de Weygaert, 2024; Kugel & van de Weygaert, 2024; Libeskind et al., 2018). The same tidal forces also transfer AM to halos and galaxies during collapse.
@@
While TTT has been successful in explaining the initial acquisition of AM and its early alignment with the LSS in a statistical sense, several critical aspects remain poorly understood. In this work, we focus on unresolved questions, such as the role of different tidal field scales in AM acquisition, the connection between these scales and protohalo properties, and the timing and nature of deviations from TTT predictions during protohalo evolution. More broadly, we ask: who are the true spinning masters? That is, which factors are most influential in shaping the angular momentum acquired by DM halos: the spatial distribution of particles within the protohalo, the surrounding tidal field, the scale at which the latter is defined, or the time evolution of these quantities? By addressing these questions, we aim to advance our understanding of AM evolution, its connection to present-day halo properties and, if possible, to observable galaxy properties.
@@
Numerical simulations have played a crucial role in validating and refining TTT. While the theory accurately describes AM growth in the linear regime (White, 1984), simulations reveal significant deviations during late time evolution (Sugerman et al., 2000; Lee & Pen, 2000; Porciani et al., 2002a, b). These discrepancies emerge as environmental effects, mergers, and vortical flows become important (Vitvitska et al., 2002; Libeskind et al., 2013; Laigle et al., 2015; Bett & Frenk, 2016; López et al., 2019; Lee & Moon, 2022). Such results highlight the need to extend TTT with additional physics that accounts for non-linear [^1] Hereafter we use the term “non-linear” in the loose sense common in the literature, referring to the late stages of halo evolution when the predictions of linear theory are no longer adequate to describe the dynamics of the matter. We also use it to denote processes typical of this stage, such as mergers, shell crossing, and fly-bys. In the very late phase ($z<1$), we use the term “highly non-linear” simply to indicate a regime where the linear approximation fails at any scale relevant for the halos in our study. In the same sense, the term “quasi-linear” is used to refer to intermediate redshifts ($z\sim 20-4$), when we see the emergence of the first non-primordial structural patterns. processes and environmental factors (e.g. Codis et al., 2015; Wang & Kang, 2017; Neyrinck et al., 2020; Ebrahimian & Abolhasani, 2021; Moon & Lee, 2023; López & Merchán, 2023).
@@
First, while TTT provides an elegant framework, simulations reveal significant deviations in DM halo AM evolution from its predictions. This suggests either theoretical incompleteness or gaps in our understanding of certain aspects, particularly the coupling scale between tidal fields and protohalo inertia tensors in Lagrangian space. While studies show that varying density field smoothing scales affect collapse times and inertia-tidal alignments (Lee et al., 2009; Hahn et al., 2009; Moon & Lee, 2024), the conventional assumption that protohalo size determines the relevant TTT scale lacks rigorous validation. Critical unanswered questions remain: How do tidal field properties systematically vary with filtering scale? Which scales dominate tidal-inertia tensor correlations? How are these scales linked to protohalo characteristics? A thorough analysis of these scale-dependent effects is notably absent in current literature.
@@ 2 Brief outline of the TTT
Figure 1: Five examples of the evolution of protohalos, with increasing mass from top to bottom. Each column, from left to right, shows the particle distribution at $z=80,4.5,1.5,0.5$ and 0, respectively. The black dots represent the full FoF group, while the red dots show the particles that reside within $0.1R_{\mathrm{vir}}$ at present time (see Section 6). Notice that the axes that correspond to each protohalo are scaled differently.
This equation indicates that, during the early stages of structure formation, the AM arises from the misalignment between the shape of the protohalo and the principal directions of expansion or collapse associated to its surrounding tidal field appropriately smoothed. The ambiguity of this expression is deliberate: in general, it is not prescribed what is the scale of the tidal field that couples with the inertia tensor to produce AM. It is usually assumed that corresponds roughly to the size of each protohalo. However, (1) in Lagrangian space, protohalo boundaries are highly irregular and may be non–simply-connected (see Figure 1), so their size and shape can only be established a posteriori from the particle membership assigned by a halo finder; moreover, while such criteria are well defined in Eulerian space at the time of identification, their mapping to the Lagrangian region is less direct and introduces additional arbitrariness, and (2) the amplitude and orientation of the tidal field can change with the smoothing scale both below and above the characteristic size of each protohalo. Hence, even with an unequivocal definition of the protohalo size $R_{\mathrm{h}}$, the resulting tidal field, and thus the TTT predictions, can be highly sensitive to the smoothing scale, differing substantially when the density field is smoothed over, e.g., $R_{\mathrm{h}}/2$, $R_{\mathrm{h}}$, or $2R_{\mathrm{h}}$.
@@ 3.1 The simulation, halos and protohalos
In this work we use the code gadget 2 (Springel, 2005) to simulate the cosmological evolution of $1600^{3}$ dark matter particles in a periodic cube of side $400{\textrm{ $h^{-1}$Mpc}}$, with a mass per particle $m_{\mathrm{p}}=1.18219\times 10^{9}{\textrm{ $h^{-1}$M${}_{\sun}$}}$. This simulation consists of 205 outputs between $z=80$, when the initial conditions were generated, and $z=0$, which corresponds to the present time. Cosmological parameters were taken from the Planck Collaboration results (Planck Collaboration et al., 2020): matter density $\Omega_{\mathrm{m}}=1-\Omega_{\Lambda}=0.315$, Hubble constant $H_{0}=67.4{\textrm{ km\,s${}^{-1}$}}{\textrm{ Mpc}}^{-1}$ and normalization parameter $\sigma_{8}=0.811$. Dark matter halo identification was performed at $z=0$ using a standard Friends of Friends (FoF) algorithm with percolation length $l=0.17\bar{\nu}^{-1/3}$, where $\bar{\nu}$ is the mean particle number density. In order to avoid biases in the determination of dynamical properties due to the low number of particles (see e.g. Paz et al., 2006; Bett et al., 2007), our analysis was carried out discarding FoF groups with less than $250$ particles. The result is a sample of about $6.5\times 10^{5}$ halos with masses $M_{\mathrm{FoF}}\geqslant 3\times 10^{11}{\textrm{ $h^{-1}$M${}_{\sun}$}}$.
@@
In addition to analyzing the full FoF groups, we track the particles that, by $z=0$, reside within fixed fractions of the virial radius ($0.1R_{\mathrm{vir}}$ and $0.5R_{\mathrm{vir}}$) of each halo (see Section 6). By comparing these inner particles to their outer counterparts, we aim to understand whether the AM assembly is driven differentially (e.g., if inner particles experience distinct tidal torques or non-linear effects) or if the entire Lagrangian region evolves coherently. This distinction is important for understanding how the AM of the final halo relates to its initial conditions.
Figure 1 presents the Lagrangian evolution of five representative FoF groups (black dots) and their inner $0.1R_{\mathrm{vir}}$ particles (red dots) at redshifts $z=80,4.5,1.5,0.5$, and 0. The full FoF groups exhibit diverse initial configurations, ranging from contiguous to fragmented spatial distributions, particularly for higher-mass halos. This morphological variety demonstrates the complexity of characterizing Lagrangian regions. The inner particles show similarly irregular distributions, either scattered throughout the entire volume or partially clustered in subregions.
The evolutionary sequence reveals distinct phases of collapse. By $z=4.5$, particles remain broadly distributed, with only faint filamentary structures emerging (a pattern reminiscent of large-scale structure formation; see Neyrinck et al. 2025 for a detailed analysis of such “halo root systems”). The distribution becomes increasingly clumpy by $z=1.5$, with most inner particles forming an offset core that becomes prominent by $z=0.5$.
@@
Finally, we need to characterize the large-scale structure at present time. To this end, we use the code NEXUS+ (Cautun et al., 2013) at $z=0$, which establishes the type of region in which each halo is embedded through an analysis of the Hessian of the smoothed local density field at different scales. In this way, it identifies not only the cosmological environment (void, wall, filament or node), but also the preferred directions of collapse and expansion of matter, $\hat{\mathbf{e}}_{k}$, with $k=1,2,3$. In the case of filamentary regions, for example, $\hat{\mathbf{e}}_{3}$ can be associated with the direction of the main axis or spine. A more detailed description of NEXUS+ can be found on Appendix A.
@@ 3.1.1 The size of protohalos
Our second measure of protohalo size is the equivalent radius, which we define as the radius of a sphere that is equal in volume to the ellipsoid associated with the eigenvalues of the inertia tensor. Hence, given the semiaxes $a,b,c=\sqrt{\mathrm{i}_{k}}$ for $k=1,2,3$, the equality $\frac{4}{3}\pi R_{\mathrm{eq}}^{3}=\frac{4}{3}\pi\,abc$ allows to compute the equivalent radius as:
$R_{\mathrm{eq}}=(abc)^{1/3}.$
@@
This quantity is computed using the particle distribution of each protohalo at $z=80$, i.e., at the same time when we analyze its surrounding tidal field.
@@
Figure 2: Protohalo size as a function of their mass. The dot, solid and dashed lines show, respectively, the median of the equivalent radius ($R_{\mathrm{eq}}$), the Lagrangian radius ($R_{\mathrm{lag}}$), and ten times the virial radius ($10R_{\mathrm{vir}}$) for protohalos of different mass. While the value of $R_{\mathrm{lag}}$ is computed exactly from the halo mass, $R_{\mathrm{eq}}$ and $10R_{\mathrm{vir}}$ present some scatter which can be seen as the dispersion of the gray circles and crosses. The colored regions indicate the three mass bins we use in our analysis.
Finally, we have $10R_{\mathrm{vir}}$, i.e. ten times the virial radius of each present-day halo. While $R_{\mathrm{vir}}$ is measured at $z=0$, $10R_{\mathrm{vir}}$ is treated as a comoving size that serves as a fixed scale at any redshift. The factor multiplying $R_{\mathrm{vir}}$ aims to take into account that, at early times, the protohalo particle distribution is significantly more extended in comoving coordinates than the corresponding virialized object at $z=0$. The value of this factor has been chosen such that $10R_{\mathrm{vir}}$ represents a scale typically larger than $R_{\mathrm{lag}}$ at $z=80$. Moon & Lee (2024) found that calculating the tidal field using a similar filter length, $8R_{\mathrm{vir}}$, yields a universal correlation with the inertia tensor, independent of the mass of each protohalo. Moreover, they found that $8R_{\mathrm{vir}}$ establishes an upper limit beyond which the so-called “type II spin transition” is not observed, i.e., the transition of $\hat{\mathbf{J}}$ from $\perp\hat{\mathbf{t}}_{1}$ to $\parallel\hat{\mathbf{t}}_{1}$, where $\hat{\mathbf{t}}_{1}$ represents the first axis of the tidal field. Thus, the value $10R_{\mathrm{vir}}$ allows us to test these results and compare our conclusions.
Figure 2 shows the relation between halo mass and the three different measures of the protohalo size at $z=80$: the equivalent radius ($R_{\mathrm{eq}}$), the Lagrangian radius ($R_{\mathrm{lag}}$) and ten times the virial radius ($10R_{\mathrm{vir}}$). Notice that, whereas the value of $R_{\mathrm{lag}}$ is computed exactly from the halo mass, $R_{\mathrm{eq}}$ and $10R_{\mathrm{vir}}$ present some scatter which can be seen as the dispersion of the gray circles and crosses, respectively. Naturally, the protohalo size tends to increase with halo mass. The equivalent radius spans a range of values such that its median goes from $\sim 0.5$ to $2.6{\textrm{ $h^{-1}$Mpc}}$ in the mass range $3\times 10^{11}-5\times 10^{13}]{\textrm{ $h^{-1}$M${}_{\sun}$}}$. Typically, for each protohalo, $R_{\mathrm{eq}}$ is about half of the value of $R_{\mathrm{lag}}$ and $\sim 1/3$ of $10R_{\mathrm{vir}}$.
@@
In order to consider the possible dependence of our results with halo mass, we split our halo population in three subsamples, shown in Figure 2 as coloured regions. With [^2] Hereafter, we use the convention $\log\equiv\log_{10}$. $m=\log\left[M_{\mathrm{FoF}}/({\textrm{ $h^{-1}$M${}_{\sun}$}})\right]$, we define a low mass subsample within the range $m=[11.5,12.0]$, a subsample of intermediate mass $m=[12.0,12.7]$ and, finally, a high mass subsample with masses in the range $m=[12.7,13.7]$. These ranges have been chosen such that a well known trend in spin alignment at $z=0$ can be easily identified: we expect the spin of the low (high) mass subsample to be preferentially aligned (perpendicular) to the present day filaments.